OpenStudy (gio333):

A line has a slope of –5 and a y-intercept of (0, 3). What is the equation of the line that is perpendicular to the first line and passes through the point (3, 2)?

OpenStudy (nincompoop):

what does it mean when a line is perpendicular to another line?

OpenStudy (anonymous):

Find the slope of the perpendicular line by doing m_1 * m_2 =-1

OpenStudy (anonymous):

Perpendicular equations are the opposite signs of the given points. at least that's how I learned how to do it.

OpenStudy (anonymous):

y=mx+b

OpenStudy (nincompoop):

a line is perpendicular to another line IF its slope is a negative reciprocal of the other line example: if the first line has a slope of 5 then the line perpendicular to it has a slope of - 1/5 if the first line has a slope of -2/5 then the line perpendicular to it is 5/2 do you need further clarification?

OpenStudy (nincompoop):

@Gio333 if you are not going to respond, you might as well log off and go to sleep

OpenStudy (gio333):

sorry i was trying to figure it out

OpenStudy (dan815):

y- 2=1/5(x-3,)

OpenStudy (gio333):

is the answer y=-2x+7?

OpenStudy (dan815):

if m1=a/b and m2 is perpendiculap slope then = -b/a

OpenStudy (dan815):

a line that passes through a point (a,b) and has slope m= y-b=m(x-a)

OpenStudy (dan815):

you can re arrange this equation for a slope intercept form y=mx-ma+b

OpenStudy (nincompoop):

do you still need help?